OpenStudy (anonymous):

Prove: sec^2x + tan^2x = (1-sin^4x)sec^4x

hartnn (hartnn):

what have you tried ?

hartnn (hartnn):

start from right... (1-sin^4x)sec^4x = sec^2x (sec^2x - sec^2x sin^4x)

OpenStudy (anonymous):

whats next ??

hartnn (hartnn):

sec^2x (sec^2x - sec^2x sin^4x) = sec^2x (1+tan^2x - (1/cos^2x) sin^4x)

hartnn (hartnn):

getting too complicated ??

hartnn (hartnn):

then convert everything into sin and cos...mostly works...

OpenStudy (anonymous):

yes the power 4 is confusing

OpenStudy (anonymous):

the next step is l.c.m ??

hartnn (hartnn):

would you like me to continue or convert everything to sin cos ?

OpenStudy (anonymous):

no , not sin cos continue the way it is

hartnn (hartnn):

sec^2x (1+tan^2x - (1/cos^2x) sin^4x) = sec^2x (1+tan^2x - (sin^2 x/cos^2x) sin^2x)

hartnn (hartnn):

i need to find an easier way...

hartnn (hartnn):

got it

OpenStudy (anonymous):

:)

hartnn (hartnn):

(1-sin^4x)sec^4x = (sec^4x - sin^4 x sec^4 x) = (sec^4 x - tan^4 x)

hartnn (hartnn):

=(sec^2 x+ tan^2 x)(sec^2x - tan^2x) ans whats sec^2x - tan^2x =..?

hartnn (hartnn):

identities used are \(\sin x*\sec x= \sin x/\cos x = \tan x\) and \(a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

1

hartnn (hartnn):

which step you need explaning ? :P

OpenStudy (anonymous):

nothing i got it, it was easy

hartnn (hartnn):

glad to hear :)

OpenStudy (anonymous):

\[\frac{ \cos^3x-\sin^3x }{ cosx-sinx }=1+sinx * cosx\]

OpenStudy (anonymous):

^ can u tell me start of this prove

hartnn (hartnn):

\(\large (a^3-b^3)=(a-b)(a^2+ab+b^2)\)

hartnn (hartnn):

@koli123able

hartnn (hartnn):

and \(\sin^2x+\cos^2x=1\)

OpenStudy (anonymous):

thanks u r the best !!!

hartnn (hartnn):

not really, but i'll take the compliment :P :)